OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 4 mm and a height of 7.2 mm

OpenStudy (saifoo.khan):

Use the formula.

OpenStudy (saifoo.khan):

\[Volume = \text{base area} \times height\]

OpenStudy (saifoo.khan):

\[volume = \pi r^2 h\]

OpenStudy (saifoo.khan):

\[Volume = \pi (2)^2 (7.2)\]

OpenStudy (saifoo.khan):

90.478 units cube.